OpenStudy (anonymous):

HOW DO WRITE LINEAR EQUATIONS IN STANDARD FORM??? 10y-3x+6=11

OpenStudy (anonymous):

oh no i see you have a \(+6\) on the left and an \(11\) on the right subtract

OpenStudy (mathstudent55):

\(ax + by + c = 0 \)

OpenStudy (anonymous):

wat do i subtract

OpenStudy (anonymous):

subtract 11 on both sides

OpenStudy (anonymous):

so i subtract 11 from 6 or or 10

OpenStudy (anonymous):

from the 6

OpenStudy (anonymous):

\[10y-3x+6=11\] \[10y-3x=5\] \[3x-10y=-5\] is probably what they have in mind

OpenStudy (anonymous):

^^ i thnik hes write i accidently gave you the quadratic from XD i apologize

OpenStudy (anonymous):

why did you switch the 10y and 3x and make the 5 a negative

OpenStudy (anonymous):

????

OpenStudy (anonymous):

you can always ya know....use google too to see what the different forms are and do the equation accordingly...

OpenStudy (anonymous):

ok v.v

OpenStudy (anonymous):

im sorry =( or you can try one of their answers

OpenStudy (agent0smith):

Standard form is Ax+By=C, A, B, C are all integers, GCF of 1, and A is positive. So, get all the x's and y's on one side, get all the numbers on the other side.